OpenStudy (anonymous):

Find the mean for the given set of data. 18, 15, 22, 18, 25, 19, 23

OpenStudy (anonymous):

no i will not tell u but what i will tell u is this: the arithmetic mean of these numbers is greater than geo mean kk?

OpenStudy (anonymous):

the mea is fgenerally the average so (18+15+22+18+25+19+23)=140/7= 20. the mean is 20

OpenStudy (amistre64):

spose we subtract 15 from all these 15 + (3, 0, 7, 3, 10, 4, 8) the mean is the same as if we take the mean of the right part and add 15 back to it 15 + (35/7) 15 + (5) = 20